OpenStudy (anonymous):

State the domain of the rational function. f(x) =7/14-x

OpenStudy (s3a):

negative infinity to infinitely closer to 14 and then from infinitely close to 14 to positive infinity. To write it out, x ∈ ]-∞,14[ ∪ ]14,∞[ or, equivalently, x ∈ (-∞,14) ∪ (14,∞)

OpenStudy (s3a):

The reason why 14 is the number of choice is because the denominator of the fraction is 0 when x = 14 since 14 - x = 14 - 14 = 0 (when x = 14). This is also why the "legal" values approach 14 from the 13.9999999999... and 14.0000000000000...1 sides (since 14 cannot be reached since all hell breaks loose if you divide by 0).

OpenStudy (anonymous):

Thanks, that helped alot:)

OpenStudy (s3a):

No problem. :)